OpenStudy (anonymous):

plz help urgent!!! Show all work in dividing (y^2-y-12)/(y^2-6y+8 by (y^2+2y-15)/(y^2-7y+10) Use complete sentences to explain how to simplify this expression.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@nicjohns

OpenStudy (anonymous):

@BlackLai

OpenStudy (anonymous):

@dumbsearch2

OpenStudy (dumbcow):

factor everything....then multiply by reciprocal of denominator ex: \[\large \frac{\frac{2}{x}}{\frac{3}{y}} = \frac{2}{x}*\frac{y}{3}\]